# EPN: An Ego Vehicle Planning-Informed Network for Target Trajectory Prediction

### v1

@@ Abstract
Trajectory prediction plays a crucial role in improving the safety and reliability of autonomous vehicles, serving as an intermediate link between perception and planning. However, due to the highly dynamic and multimodal nature of the task, accurately predicting the future trajectory of a target vehicle remains a significant challenge. To address these challenges, we propose an Ego vehicle Planning-informed Network (EPN) for multimodal trajectory prediction. Current trajectory prediction methods typically use the historical trajectory and vehicle attributes as inputs, focusing primarily on how historical information influences the future trajectory of the target vehicle. In real-world driving scenarios, however, the future trajectory of a vehicle is influenced not only by its own historical data but also by the behavior of other vehicles on the road. To address this, we incorporate the future planned trajectory of the ego vehicle as an additional input to simulate the mutual influence between the ego vehicle’s planned trajectory and the predicted trajectory of the target vehicle. Furthermore, to tackle the challenges of intention ambiguity and large prediction errors often encountered in methods based on driving intentions, we propose a target’s endpoint prediction module. This module first predicts the possible endpoints of the target vehicle, then refines these predictions through a correction mechanism, and finally generates a complete multimodal predicted trajectory based on the corrected endpoints. Experimental results demonstrate that, compared to other trajectory prediction methods, EPN achieves an average reduction of 34.9%, 30.7%, and 30.4% in RMSE, ADE, and FDE evaluation metrics on the NGSIM dataset, and an average reduction of 64.6%, 64.5%, and 64.3% in RMSE, ADE, and FDE on the HighD dataset. These results highlight the strong performance of EPN in trajectory prediction. The code is available at https://github.com/whutpsq/EPN.
Index Terms:
Autonomous driving, trajectory prediction, ego vehicle planning, endpoint correction, LSTM, CVAE.
@@ I Introduction
With the rapid development of autonomous driving technology, the safety of autonomous vehicles has garnered increasing attention. Accurately predicting the future trajectory of surrounding vehicles is crucial for the safe operation of autonomous vehicles[2]. As an upstream module of planning, trajectory prediction plays a key role in supporting planning tasks and is one of the critical components in enhancing the safety of autonomous driving systems. However, due to the high dynamic and multimodal nature of trajectory prediction tasks, accurately forecasting the future trajectory of a target vehicle remains a significant challenge. High dynamism refers to the influence of other traffic participants on the target vehicle, which may cause the target vehicle’s trajectory to change unexpectedly. Multimodality refers to the possibility of multiple plausible future trajectories for a given historical trajectory. These characteristics make trajectory prediction a complex and difficult task.
In the early stages of trajectory prediction technology, methods based on physical models were commonly used. Dynamic models made short-term predictions by describing the vehicle’s dynamic characteristics[3][4]. However, these models are often complex, and their performance improvement in trajectory prediction is limited despite the increased model complexity. Kinematic models, on the other hand, are simpler, primarily using kinematic attributes such as vehicle speed, acceleration, and steering angle[5][6]. However, these models are only suitable for relatively simple scenarios. These traditional methods primarily focus on the state of the target vehicle and fail to account for the influence of other factors, such as surrounding vehicles and environmental conditions. The Kalman filter method incorporates noise consideration within the model construction and predicts vehicle trajectories through Gaussian distribution modeling.[7][8]. Compared to the Kalman filter, which is limited to linear scenarios, the Monte Carlo method offers advantages in handling complex nonlinear and multimodal scenarios[9]. However, practical applications require a balance between efficiency and accuracy.
As technology has advanced, classic machine learning methods have been increasingly applied to vehicle trajectory prediction. Unlike physics-based models, machine learning methods are data-driven approaches. Notable methods include Gaussian Processes (GP), Support Vector Machines (SVM), Hidden Markov Models (HMM), and Dynamic Bayesian Networks (DBN). GP use historical trajectory data to learn the potential distribution of future trajectories[10]. SVM classify data by finding an optimal hyperplane, helping to determine driving intentions such as left turns, straight movements, and right turns[11]. HMM effectively simulate temporal changes and uncertainties in driving behavior by modeling vehicle driving patterns as hidden states, but they are limited to systems with discrete states[12][13]. DBN are well-suited for handling complex multidimensional states and long-term dependencies, but they can suffer from significant errors when converting recognized intentions into accurate trajectories[14][15].
With the rapid development of deep learning, neural network models have made significant advancements in vehicle trajectory prediction, encompassing Recurrent Neural Network (RNN), Convolutional Neural Network (CNN), Graph Neural Network (GNN), Generative Adversarial Network (GAN) and Conditional Variational Auto Encoder (CVAE)[2][16]. Researchers often leverage one or more neural network architectures to achieve better prediction accuracy.
Owing to their extensive parameterization, neural networks can effectively model not only vehicle dynamics and attributes but also complex interactions among traffic participants and road environment features.
Deo et al.[17] proposed a convolutional social pooling module based on [18], which extracts the interaction relationships between vehicles at a deeper level. Ding et al.[19] introduced a vehicle behavior interaction network using recurrent neural networks to model interactions based on the hidden states of vehicles. Xin et al.[20] developed a trajectory prediction method utilizing two LSTMs, which first identifies the driver’s intention and then predicts the future trajectory of the vehicle based on that intention. Gupta et al.[21] applied generative adversarial networks to predict the trajectories of target vehicles. Wang et al.[22] proposed a social pooling method inspired by wave superposition to dynamically model both local and global interactions.
Incorporating prior knowledge significantly enhances trajectory prediction accuracy. For instance, high-definition maps have been widely used to improve predictions. Casas et al.[23] jointly tackled vehicle detection and trajectory prediction by processing raw nt clouds and gridded high-definition maps. Gao et al.[24] introduced vectorized encoding for high-definition maps and traffic participants, reducing information loss typically caused by rasterization, lowering computational complexity, and improving prediction accuracy. Subsequent studies, such as [25] -[28], extensively utilized vectorized encoding methods.
Beyond fully utilizing available information, researchers also aim to improve prediction accuracy by incorporating additional data. Song et al.[29] were among the first to consider how different ego vehicle plans impact the future trajectory of a target vehicle. They used an LSTM encoder to model the ego vehicle’s planned trajectory and extracted interaction features through a convolutional social pooling module. Their experiments demonstrated that incorporating ego vehicle planning information significantly improves model prediction accuracy. Guo et al.[30] further enhanced model performance by using a spatiotemporal dual-attention mechanism to extract spatiotemporal features while incorporating autonomous vehicle planning, achieving higher trajectory prediction accuracy. Sheng et al.[31] further reduced uncertainty in the target vehicle’s behavior by explicitly encoding ego vehicle’s planned trajectory into a graph convolutional network.
The characteristics of input information significantly influence trajectory prediction models, while the output format of these models is also crucial in determining their performance. Multimodal trajectory prediction typically employs two approaches. The first is driving intention-based prediction, where driver behaviors such as left turns, right turns, straight movements, acceleration, and deceleration are classified, and trajectories are predicted for each intention, as seen in studies like [29], [30], [32] and [33]. The second is endpoint-based prediction, such as [26], [34] and [35], which first predict the target vehicle’s potential endpoint and subsequently generate the complete trajectory based on the predicted endpoint.
In this paper, we propose a multimodal trajectory prediction model based on endpoint correction and ego vehicle planning. This model takes as input the historical trajectories and state information of all vehicles in the scene, along with the planned trajectory of the ego vehicle, and outputs multimodal predictions for the future trajectory of the target vehicle. We intentionally did not incorporate high-definition maps due to their high production and updating costs. With the current trend in intelligent driving solutions moving towards lightweight maps, trajectory prediction methods that do not rely on high-definition maps still hold significant research value. Our contributions can be summarized as follows:
@@
    We introduce a feature fusion encoding module that first extracts temporal features from the input information using an LSTM encoder, then models interactions between various traffic participants with a convolutional social pooling network to capture spatial features of the scene. The module outputs a unified feature vector that integrates both temporal and spatial characteristics.
@@
    We propose a target’s endpoint prediction module that first predicts the possible endpoints of the target vehicle using a CVAE, then refines the preliminary predicted endpoints coordinates with a correction mechanism, and finally uses the corrected endpoints to predict the complete trajectory of the target vehicle.
The remainder of the paper is organized as follows: Section II presents the problem formulation, Section III describes the model architecture, Section IV details the experimental setup and results, and Section V concludes the paper with a summary.
II PROBLEM FORMULATION
@@ II-B Model Output
III MODEL ARCHITECTURE
This article introduces an Endpoint-based Planning-informed Network (EPN) for multimodal trajectory prediction by integrating vehicle driving status information with the future planning trajectory of the ego vehicle, the model architecture is shown in Figure 2. EPN consists of three main modules: the feature fusion encoding, the target’s endpoint prediction, and the LSTM decoder. The feature fusion encoding module separately encodes the information from the ego vehicle, target vehicle, and adjacent vehicles. The encoded features of the ego vehicle and adjacent vehicles are processed through a convolutional social pooling network to extract interaction information, generating social feature vectors. Simultaneously, the target vehicle’s encoded features are transformed into dynamic feature vectors via a fully connected layer. These vectors are concatenated to form environmental feature vectors. The target’s endpoint prediction module employs CVAE to predict potential target vehicle endpoints and refines them to improve accuracy. Finally, the LSTM decoder module combines the environmental and endpoint feature vectors to produce the target vehicle’s complete multimodal predicted trajectory.
@@ III MODEL ARCHITECTURE
III-A Feature Fusion Encoding
@@
$e_{dyn}=FC(e_{tar}^{t})$ (12)
here, $e_{nei}^{t}$ represents the final encoded feature vector of the vehicles surrounding the target vehicle at time $t$, which includes encoded information about the ego vehicle, while $e_{tar}^{t}$ is the final encoded feature vector of the target vehicle at time $t$. $e_{dyn}$ denotes the dynamic feature vector of the target vehicle after dimensionality reduction.
@@
To effectively model the interaction between the ego vehicle, adjacent vehicles, and the target vehicle—and to emphasize the influence of adjacent vehicles’ historical information and the ego vehicle’s planned future trajectory on the target vehicle’s future path—we introduce a convolutional social pooling network to simulate inter-vehicle interactions within the scene. First, a mesh centered on the target vehicle is established using a masking mechanism, which collects the feature vectors of surrounding vehicles into a social tensor. The masking calculation formula is as follows:
@@
here $gri{d_{i,j}}$ indicates whether an adjacent vehicle occupies position $\left({i,j}\right)$ in the grid centered on the target vehicle. If $gri{d_{i,j}=1}$, it signifies the presence of an adjacent vehicle at that position, and $mas{k_{i,j}}$ is set to 1; otherwise, $mas{k_{i,j}}$ is set to 0. The resulting social tensor after masking adjacent vehicles is denoted as $tenso{r_{nei}}$, while the social tensor for the planned trajectory of the ego vehicle is represented as $tenso{r_{plan}}$.
TABLE I: Structure and Parameters of Convolutional Social Pooling Network
Layer Name Function & Parameter
Conv1 Conv2d(64,64,3)
Relu LeakyRelu(0.1)
MP1 MaxPool2d((3,3),2)
Conv2 Conv2d(64,16,(3,1))
MP2 MaxPool2d((2,2),(1,0))
Figure 3: The Structure of Convolutional Social Pooling Network.
The social tensor is subsequently input into the convolutional pooling network to extract interaction information between vehicles. Table I presents the structure and parameters of the convolutional social pooling network, which consists of two convolutional layers, two pooling layers, and a ReLU layer. Figure 3 illustrates the tensor transformation process within the convolutional social pooling network. In this architecture, each convolutional layer is followed by a ReLU layer, employing $LeakyRelu(0.1)$ as the activation function. The primary role of the convolutional layers is to extract interactive information from the social tensors through convolutional operations, while the pooling layers downsample features to reduce computational complexity and enhance the model’s translational invariance. For the social tensors representing adjacent vehicles and the ego vehicle planning, interaction information is extracted using the ”convolution-pooling-convolution” operation, resulting in the intermediate tensors $en{c_{nei}}$ and $en{c_{plan}}$. These tensors are then concatenated through a pooling layer to generate the social feature vector, denoted as $en{c_{social}}$, which encapsulates the spatial feature information of the target vehicle. The formulas are as follows:
$en{c_{nei}}=Relu(Conv2(MP1(Relu(Conv1(tenso{r_{nei}})))))$ (15)
$en{c_{plan}}=Relu(Conv2(MP1(Relu(Conv1(tenso{r_{plan}})))))$ (16)
@@
III-B Target’s Endpoint Prediction
@@
In this module, the model extracts endpoint features using different processes at various stages. During the training phase, distinct multilayer perceptrons (MLPs) are employed as the endpoint encoder, latent variable encoder, and latent variable decoder. First, the endpoint encoder encodes the target vehicle’s true endpoint. The resulting feature vector is concatenated with the environmental feature vector and passed to the latent variable encoder to produce a latent variable. The latent variable’s mean $\mu$ and standard deviation $\sigma$ are calculated, and Gaussian noise $z$ is sampled from a normal distribution $N(\mu,\sigma)$. Finally, $z$ is concatenated with the environmental feature vector and decoded by the latent variable decoder to generate the target vehicle’s endpoint. The formulas for the training phase are provided in equations (19) to (21):
$end\_feature={E_{end}}(G)$ (19)
$latent={E_{latent}}(concat(enc,end\_feature))$ (20)
@@
where $G$ represents the true endpoint of the target vehicle, while ${E_{end}}$, ${E_{latent}}$, and ${D_{latent}}$ denote the endpoint encoder, latent variable encoder, and latent variable decoder, respectively. $end\_feature$ refers to the encoded endpoint feature vector, and $\overline{G}$ represents the predicted endpoint of the target vehicle. During the validation and testing phases, since the true endpoint of the target vehicle is unknown, we directly sample Gaussian noise $z$ from the latent space based on a normal distribution $N(0,{\sigma_{T}}I)$. The sampled noise $z$ is then concatenated with the environmental feature vector and decoded to generate potential endpoint predictions. Following the truncation trick in PECnet[36], we set the standard deviation ${\sigma_{T}=1.3}$. Additionally, by controlling the number of times the latent space is sampled, multiple possible endpoint coordinates for the target vehicle can be predicted.
@@
Due to the significant error in directly predicting the target vehicle’s endpoint, we employ a endpoint correction mechanism to improve the accuracy of the prediction. After the first-stage prediction results are generated, the correction mechanism is applied to refine the predicted endpoint. Specifically, we first use a endpoint encoder to re-encode the predicted endpoint $\overline{G}$ of the target vehicle. The resulting feature vector is concatenated with the environmental feature vector and input into a endpoint decoder ${D_{end}}$, which has the same structure as the latent variable decoder ${D_{latent}}$. This decoder generates a deviation value that represents the $offset$ between the predicted endpoint and the true endpoint. Finally, the $offset$ is added to the first-stage predicted endpoint to obtain the corrected endpoint $\widehat{G}$. The detailed calculation process is shown in equations (22) to (24):
$refine\_end\_feature={E_{end}}(\overline{G})$ (22)
$offset={D_{end}}(concat(enc,refine\_end\_feature))$ (23)
$\widehat{G}=\overline{G}+offset$ (24)
III-C LSTM Decoder

### v2

@@ Abstract
Trajectory prediction plays a crucial role in improving the safety of autonomous vehicles. However, due to the highly dynamic and multimodal nature of the task, accurately predicting the future trajectory of a target vehicle remains a significant challenge. To address this challenge, we propose an Ego vehicle Planning-informed Network (EPN) for multimodal trajectory prediction. In real-world driving, the future trajectory of a vehicle is influenced not only by its own historical trajectory, but also by the behavior of other vehicles. So, we incorporate the future planned trajectory of the ego vehicle as an additional input to simulate the mutual influence between vehicles. Furthermore, to tackle the challenges of intention ambiguity and large prediction errors often encountered in methods based on driving intentions, we propose an endpoint prediction module for the target vehicle. This module predicts the target vehicle endpoints, refines them using a correction mechanism, and generates a multimodal predicted trajectory. Experimental results demonstrate that EPN achieves an average reduction of 34.9%, 30.7%, and 30.4% in RMSE, ADE, and FDE on the NGSIM dataset, and an average reduction of 64.6%, 64.5%, and 64.3% in RMSE, ADE, and FDE on the HighD dataset. The code will be open sourced after the letter is accepted.
@@ I INTRODUCTION
With the rapid development of autonomous driving technology, the safety of autonomous vehicles has garnered increasing attention. Accurately predicting the future trajectory of surrounding vehicles is crucial for the safe operation of autonomous vehicles [1]. As an upstream module of planning, trajectory prediction plays a key role in supporting planning tasks and is one of the critical components in enhancing the safety of autonomous driving systems. However, due to the high dynamic and multimodal nature of trajectory prediction tasks, accurately forecasting the future trajectory of a target vehicle remains a significant challenge. High dynamism refers to the influence of other traffic participants on the target vehicle, which can cause the trajectory of the target vehicle to change unexpectedly. Multimodality refers to the possibility of multiple plausible future trajectories for a given historical trajectory. These characteristics make trajectory prediction a complex and difficult task.
To address these challenges, we propose an Ego vehicle Planning-informed Network (EPN) for multimodal trajectory prediction. In response to the high dynamic characteristics of trajectory prediction, current trajectory prediction methods typically use the historical trajectory and vehicle attributes as inputs, focusing primarily on how historical information influences the future trajectory of the target vehicle. However, in real-world driving scenarios, the future trajectory of a vehicle is influenced not only by its own historical data, but also by the behavior of other vehicles on the road. So, we incorporate the future planned trajectory of the ego vehicle as an additional input to model the mutual influence between vehicles. Furthermore, to account for the multimodal nature of trajectory prediction, we introduce a target endpoint prediction module that predicts multiple plausible target endpoints, thereby generating more realistic and accurate trajectory predictions for the target vehicle. Our contributions can be summarized as follows:
@@
    We introduce a feature fusion encoding module that first extracts temporal features from the input information using a Long Short-Term Memory (LSTM) encoder, then models interactions between various traffic participants with a convolutional social pooling network to capture spatial features. The module outputs a unified feature vector that integrates both temporal and spatial characteristics.
@@
    We propose a target’s endpoint prediction module that first predicts the possible endpoints of the target vehicle using a Conditional Variational Autoencoder (CVAE), then refines the preliminary predicted endpoints coordinates with a correction mechanism, and finally uses the corrected endpoints to predict the complete trajectory of the target vehicle.
II RELATED WORK
II-A Physics-based methods
In the early stages of trajectory prediction technology, single trajectory methods were commonly used. Dynamic models made short-term predictions by describing the vehicle’s dynamic characteristics [2], [3]. Kinematic models are mainly based on kinematic attributes such as vehicle speed, acceleration, and steering angle [4], [5]. These methods primarily focus on the state of the target vehicle and fail to account for the influence of other factors, such as surrounding vehicles and environmental conditions. The Kalman filter method incorporates noise consideration within the model construction and predicts vehicle trajectories through Gaussian distribution modeling [6], [7]. Compared to the Kalman filter, which is limited to linear scenarios, the Monte Carlo method offers advantages in handling complex nonlinear and multimodal scenarios [8].
II-B Machine learning methods
As technology has advanced, machine learning methods have been increasingly applied to vehicle trajectory prediction. Unlike physics-based models, machine learning methods are data-driven approaches. Notable methods include Gaussian Processes (GPs), Support Vector Machines (SVMs), Hidden Markov Models (HMMs), and Dynamic Bayesian Networks (DBNs). GPs use historical trajectory data to learn the potential distribution of future trajectories [9]. SVMs classify data by finding an optimal hyperplane, helping to determine driving intentions such as left turns, straight movements, and right turns [10]. HMMs effectively simulate temporal changes and uncertainties in driving behavior by modeling vehicle driving patterns as hidden states, but they are limited to systems with discrete states [11]. DBNs are well-suited for handling complex multidimensional states and long-term dependencies, but they can suffer from significant errors when converting recognized intentions into accurate trajectories [12].
II-C Deep learning methods
Deep learning has revolutionized trajectory prediction by enabling comprehensive modeling of vehicle dynamics, social interactions, and environmental constraints through neural networks’ parametric capacity [13], [14], [15], [16], [17], [18], [19], [20], [21], [22], [23], [24], [25], [26], [27]. Early works established interaction modeling frameworks using convolutional social pooling [14] and GANs [15], [16], while [17] pioneered intention-aware prediction through dual-LSTM architectures. Recent advances emphasize the critical role of additional input information. Song et al. [19] and Guo et al. [20] demonstrated significant accuracy gains by encoding ego vehicle plans via LSTM and attention mechanisms, with Sheng et al. [21] further reducing behavioral uncertainty through explicit trajectory encoding in graph networks. Another important aspect that influences performance is the output formulation of prediction models, where current multimodal approaches diverge into two distinct paradigms. Intention-based methods, such as those in [19], [20], [23], and [24], classify discrete driving maneuvers to generate corresponding trajectories. In contrast, endpoint-driven approaches, such as those in [25], [26], and [27], demonstrate superior performance through endpoint-constrained trajectory generation.
III PROBLEM FORMULATION
@@ III-B Model Output
IV MODEL ARCHITECTURE
This article introduces an Ego vehicle Planning-informed Network (EPN) for multimodal trajectory prediction, the model architecture is shown in Figure 2. EPN consists of three main modules: the feature fusion encoding, the target’s endpoint prediction, and the LSTM decoder. The feature fusion encoding module separately encodes the information from the ego vehicle, target vehicle, and adjacent vehicles. The encoded features of the ego vehicle and adjacent vehicles are processed through a convolutional social pooling network to extract interaction information, generating social feature vectors. Simultaneously, the target vehicle’s encoded features are transformed into dynamic feature vectors via a fully connected layer. These vectors are concatenated to form environmental feature vectors. The target’s endpoint prediction module employs CVAE to predict potential target vehicle endpoints and refines them to improve accuracy. Finally, the LSTM decoder module combines the environmental and endpoint feature vectors to produce the target vehicle’s complete multimodal predicted trajectory.
@@ IV MODEL ARCHITECTURE
IV-A Feature Fusion Encoding
@@
$enc_{dyn}=FC(e_{tar}^{t})$ (12)
here, $e_{nei}^{t}$ represents the final encoded feature vector of the vehicles surrounding the target vehicle at time $t$, which includes encoded information about the ego vehicle, while $e_{tar}^{t}$ is the final encoded feature vector of the target vehicle at time $t$. $enc_{dyn}$ denotes the dynamic feature vector of the target vehicle after dimensionality reduction.
@@
To effectively model the interaction between the ego vehicle, adjacent vehicles, and the target vehicle while highlighting the influence of adjacent vehicles’ historical information and the ego vehicle’s planned future trajectory on the target vehicle’s future path, we introduce a convolutional social pooling network to simulate inter-vehicle interactions within the scene. First, a mesh centered on the target vehicle is established using a masking mechanism, which collects the feature vectors of surrounding vehicles into a social tensor. The masking calculation formula is as follows:
@@
here $gri{d_{i,j}}$ indicates whether an adjacent vehicle occupies position $\left({i,j}\right)$ in the grid centered on the target vehicle. If $gri{d_{i,j}=1}$, it signifies the presence of an adjacent vehicle at that position, and $mas{k_{i,j}}$ is set to 1; otherwise, $mas{k_{i,j}}$ is set to 0. The resulting social tensor after masking adjacent vehicles is denoted as $t_{nei}$, while the social tensor for the planned trajectory of the ego vehicle is represented as $t_{plan}$.
The social tensor is subsequently input into the convolutional pooling network to extract interaction information between vehicles. For the social tensors representing adjacent vehicles and the ego vehicle planning, interaction information is extracted using the ”convolution-pooling-convolution” operation, resulting in the intermediate tensors $en{c_{nei}}$ and $en{c_{plan}}$. These tensors are then concatenated through a pooling layer to generate the social feature vector, denoted as $en{c_{social}}$, which encapsulates the spatial feature information of the target vehicle. The formulas are as follows:
$en{c_{nei}}=ReLu(Conv2(MP1(ReLu(Conv1({t_{nei}})))))$ (15)
$en{c_{plan}}=ReLu(Conv2(MP1(ReLu(Conv1({t_{plan}})))))$ (16)
@@
IV-B Target’s Endpoint Prediction
@@
In this module, the model extracts endpoint features using different processes at various stages. During the training phase, distinct multilayer perceptrons are employed as the endpoint encoder, latent variable encoder, and latent variable decoder. First, the endpoint encoder encodes the target vehicle’s true endpoint. The resulting feature vector is concatenated with the environmental feature vector and passed to the latent variable encoder to produce a $latent$ variable. The $latent$ variable’s mean $\mu$ and standard deviation $\sigma$ are calculated, and Gaussian noise $z$ is sampled from a normal distribution $N(\mu,\sigma)$. Finally, $z$ is concatenated with the environmental feature vector and decoded by the latent variable decoder to generate the target vehicle’s endpoint. The formulas for the training phase are provided in equations (19) to (21):
$en{d_{feature}}={E_{end}}(G)$ (19)
$latent={E_{latent}}(concat(enc,en{d_{feature}})$ (20)
@@
where $G$ represents the true endpoint of the target vehicle, while ${E_{end}}$, ${E_{latent}}$, and ${D_{latent}}$ denote the endpoint encoder, latent variable encoder, and latent variable decoder, respectively. $en{d_{feature}}$ refers to the encoded endpoint feature vector, and $\overline{G}$ represents the predicted endpoint of the target vehicle. During the validation and testing phases, since the true endpoint of the target vehicle is unknown, we directly sample Gaussian noise $z$ from the latent space based on a normal distribution $N(0,{\sigma_{T}}I)$. The sampled noise $z$ is then concatenated with the environmental feature vector and decoded to generate potential endpoint predictions. Following the truncation trick in PECnet [28], we set the standard deviation ${\sigma_{T}=1.3}$. Additionally, by controlling the number of times the latent space is sampled, multiple possible endpoint coordinates for the target vehicle can be predicted.
@@
Due to the significant error in directly predicting the target vehicle’s endpoint, we employ an endpoint correction mechanism to improve the accuracy of the prediction. After the first-stage prediction results are generated, the correction mechanism is applied to refine the predicted endpoint. Specifically, we first use an endpoint encoder to re-encode the predicted endpoint $\overline{G}$ of the target vehicle. The resulting feature vector $end_{refine}$ is concatenated with the environmental feature vector $enc$ and input into an endpoint decoder ${D_{end}}$. This decoder generates a deviation value $d_{offset}$ that represents the offset between the predicted endpoint and the true endpoint. Finally, $d_{offset}$ is added to the predicted endpoint in the first stage to obtain the corrected endpoint $\widehat{G}$. The detailed calculation process is shown in equations (22) to (24):
$end_{refine}={E_{end}}(\overline{G})$ (22)
$d_{offset}={D_{end}}(concat(enc,end_{refine}))$ (23)
$\widehat{G}=\overline{G}+d_{offset}$ (24)
IV-C LSTM Decoder